OpenStudy (anonymous):

Solve this system of equations. Write the answer as an ordered pair. y = x 2x + 5y = 28......and.....Solve this system of equations. Write your answer as an ordered pair. y + 6 = x 2x + 2y = 0

OpenStudy (anonymous):

for the first one i got (22,0)....and the second (0,4)

OpenStudy (anonymous):

and it said they were both wrong

OpenStudy (haseeb96):

brother here y=x now put it in equation 2 2x +5x =28 7x=28 x=4 now y=x so y=4 (x,y) = ( 4,4)

OpenStudy (anonymous):

THAT WAS MY FIRST ANSWER AND I WAS LIKE, NOPE THAT DOESNT SEEM RIGHT!!!!!!

OpenStudy (haseeb96):

y=x-6 put it in equation 2 2x +2(x-6) =0 2x + 2x -12 =0 4x =12 x=3 now put x=3 into y= x-6 y=3-6 y=-3 so (x,y) = (3,-3)

OpenStudy (haseeb96):

both are correct @kyle_bradley

OpenStudy (anonymous):

ok, thank you so much. the first one i had right but deleted it and the second one sounds good

OpenStudy (haseeb96):

in the first y=x therefore it is correct

OpenStudy (anonymous):

yeah, it makes so much more sense